# LatentQGAN: A Hybrid QGAN with Classical Convolutional Autoencoder

### v2

@@ IV-C The Autoencoder
Figure 3: FID results over training. The Fig.4 shows a visual inaccuracy and decay respectively before 350 iterations and after 700 iterations
@@ V-A Experimental details
Figure 5: FID evaluated on many models for comparison with LatentQGAN, on simulator
@@
In this study, model evaluation was performed based on the Fréchet Inception Distance (FID) metric, which is commonly used to evaluate QGANs [7, 8]. The FID metric serves as a measure of the similarity between two distributions of data by computing the Fréchet Distance between their corresponding multivariate Gaussian distributions. Mathematically, FID is calculated as follows:
$FID=\lVert\mu_{r}-\mu_{g}\rVert_{2}^{2}+Tr(\Sigma_{r}+\Sigma_{g}-2(\Sigma_{r}\Sigma_{g})^{\frac{1}{2}}),$ (10)
where $\mu_{r}$ and $\mu_{g}$ are the mean vectors, and $\Sigma_{r}$ and $\Sigma_{g}$ are the covariance matrices of the real and generated data distributions, respectively. A lower FID score indicates a higher resemblance between the generated samples and the real data distribution. While useful for gauging the similarity between generated and real data distributions, the FID does not always correspond to human evaluations [24] as shown in Fig.6. Thus, though valuable, it should be supplemented with further evaluations, including perceptual studies where human observers rate the quality of generated images.
@@ V-B Experimental results
Figure 6: Generated images with the lowest FID for LatentQGAN on the simulator. Images generated on NISQ needed 550 iterations for class 9 and 5, and 650 for class 0 to achieve minimal FID, out of 800 iterations.
Model evaluation involved a comparative analysis between generated samples and ground truth data. As we can see in Fig. 3, based on our experiments, the minimum FID was obtained after 350 to 700 iterations. As depicted in Fig.4, prior to 350 iterations, the model fails to generate visually accurate images. Subsequently, between 350 and 700 iterations, the generated images improve noticeably. However, beyond 700 iterations, there is a gradual decline in visual quality.
In the results shown in Fig. 5 and Fig. 6, LatentQGAN consistently outperforms other models like QPatchGAN and MosaiQ in terms of image quality, both visually and quantitatively, in simulations. Interestingly, LatentQGAN achieves this advantageous comparison with only five 4-qubits quantum circuits and a total of 140 parameters, while MosaiQ uses eight 5-qubits circuits and 240 parameters. Additionally, LatentQGAN requires 676 times fewer iterations than MosaiQ to achieve comparable results. Moreover, although the FID of RandomDecoder remains low, the visual outcomes shown in Fig.6 suggest an inadequate representation of the original distribution.
TABLE I: Comparison of LatentQGAN’s FID on Quantum Simulator vs. Quantum Computer. Results on the quantum computer are obtained from a single run using identical noise vectors and initial parameters as those used in simulator trainings.

### v4

@@ IV-C The Autoencoder
Figure 3: FD results over training. The Fig.4 shows a visual inaccuracy and decay respectively before 350 iterations and after 700 iterations
@@ V-A Experimental details
Figure 5: FD evaluated on many models for comparison with LatentQGAN, on simulator
@@
In this study, model evaluation was performed based on the Fréchet Distance (FD) metric, which is commonly used to evaluate QGANs [7, 8]. The FD metric serves as a measure of the similarity between two distributions of data by computing the distance between their corresponding multivariate Gaussian distributions. Mathematically, FD is calculated as follows:
$FD=\lVert\mu_{r}-\mu_{g}\rVert_{2}^{2}+Tr(\Sigma_{r}+\Sigma_{g}-2(\Sigma_{r}\Sigma_{g})^{\frac{1}{2}}),$ (10)
where $\mu_{r}$ and $\mu_{g}$ are the mean vectors, and $\Sigma_{r}$ and $\Sigma_{g}$ are the covariance matrices of the real and generated data distributions, respectively. A lower FD score indicates a higher resemblance between the generated samples and the real data distribution. While useful for gauging the similarity between generated and real data distributions, the FD does not always correspond to human evaluations as shown in Fig.6. Thus, though valuable, it should be supplemented with further evaluations, including perceptual studies where human observers rate the quality of generated images.
@@ V-B Experimental results
Figure 6: Generated images with the lowest FD for LatentQGAN on the simulator. Images generated on NISQ needed 550 iterations for class 9 and 5, and 650 for class 0 to achieve minimal FD, out of 800 iterations.
Model evaluation involved a comparative analysis between generated samples and ground truth data. As we can see in Fig. 3, based on our experiments, the minimum FD was obtained after 350 to 700 iterations. As depicted in Fig.4, prior to 350 iterations, the model fails to generate visually accurate images. Subsequently, between 350 and 700 iterations, the generated images improve noticeably. However, beyond 700 iterations, there is a gradual decline in visual quality.
In the results shown in Fig. 5 and Fig. 6, LatentQGAN consistently outperforms other models like QPatchGAN and MosaiQ in terms of image quality, both visually and quantitatively, in simulations. Interestingly, LatentQGAN achieves this advantageous comparison with only five 4-qubits quantum circuits and a total of 140 parameters, while MosaiQ uses eight 5-qubits circuits and 240 parameters. Additionally, LatentQGAN requires 676 times fewer iterations than MosaiQ to achieve comparable results. Moreover, although the FD of RandomDecoder remains low, the visual outcomes shown in Fig.6 suggest an inadequate representation of the original distribution.
TABLE I: Comparison of LatentQGAN’s FD on Quantum Simulator vs. Quantum Computer. Results on the quantum computer are obtained from a single run using identical noise vectors and initial parameters as those used in simulator trainings.